OpenStudy (anonymous):

I need this answered as soon as possible!! You can earn a metal for this!! plz help! Thanx guys! y/6 (y over 6, a fraction) = -2 Whats 'y'? a. y = -12 b. y = -3 c. y = 3 d. y = 12

OpenStudy (anonymous):

6(y/6) = 6(y + 5) y - 6y = 6y + 30 + 6y 6y ------------------ 1y = 12y + 30 -1y -1y ------------------ 0 = 11y + 30